The bill for five glasses of apple juice and four salads is $9.50, but the bill for four glasses of apple juice and five salads is $10.30. What would be the bill for a glass of juice and a salad?
Answer by jainenderkapoor(61)

Suppose one apple juice costs $x and salad costs $y
we get 5x + 4y = 9.50
4x + 5y = 10.30
Multiplying equation (1) by 5 and (2) by 4
25x + 20y = 47.50
16x + 20y = 41.20

Subtracting we get -
9x = 6.30
x = 6.30/9 = .70
Putting value of x in (1)
5(.70) + 4y = 9.50
3.5 + 4y = 9.50
4y = 9.50 - 3.50 = 6.0
y = 6.00 / 4 = 1.50
Cost of 1 glass of apple juice is $0.70 and of 1 salad is $1.50
Therefore cost of bill for 1 glass of apple juice and 1 salad is .70+1.50 = $2.20


Hope the solution is clear to you. If you have any doubt you can contact me at jainenderkapoor@gmail.com I am also available for Online Tutoring at very reasonable rates.